OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = −4,096 and a4 = 64?

OpenStudy (anonymous):

for a geometric sequence, you have to find the common ration

OpenStudy (anonymous):

Would the common ratio be -0.015625?

OpenStudy (anonymous):

-4096 = 64r^3 -4096 = 262144r r = 64?

OpenStudy (anonymous):

That's not right...

OpenStudy (anonymous):

Help, anyone?

OpenStudy (anonymous):

i think nth term is written as \[a _{n} = a*r ^{n-1}\] where r = common ratio a = first term put n=4 u have been given the value of fourth term as well as the first term

OpenStudy (anonymous):

I think the 4th term is 64, it says a4 = 64...not sure if that's the same thing

OpenStudy (anonymous):

yeah fourth term is 64

OpenStudy (anonymous):

what is the common ratio?

OpenStudy (anonymous):

first term can't be negative because we cannot determine the cube root of negative number. I feel something is wrong with the ques

OpenStudy (anonymous):

Idk...straight out of the book

OpenStudy (anonymous):

sorry cube rot can be negative

OpenStudy (anonymous):

64 = -4096 *r^(4-1)

OpenStudy (anonymous):

solve this expression u will get common ratio

OpenStudy (kmeis002):

\[ a_n = a_1r^{n-1}\] where n starts at 1 If we solve for a0 and r we can find any term. Since we know two points, its possible to solve. From the problem, we can see \[ a_1 = -4095\] Then \[a_4 = a_1*r^{4-1}= -4095(r)^{3} = 64 \] Solving for r: \[ r = \left (\frac{64}{-4095} \right )^{1/3} \] Cubed roots can be negative. So r is a negative ratio, it will be an alternating sereies. \[ r = -\frac{4}{16} = -\frac{1}{4} \] Now the sequence is defined, plug in n = 6 and solve.

OpenStudy (anonymous):

64 = -4096r^3 64 = -68719476736r r = -1073741824 That doesn't seem right...

OpenStudy (kmeis002):

You must first divide by -4096 and then take the cube root (inverse of 3rd power) to solve r.

OpenStudy (anonymous):

So -1/4 is the common ratio?

OpenStudy (anonymous):

yup -1/4 is the common ratio

OpenStudy (anonymous):

now determine sixth term using formula for nth term

OpenStudy (anonymous):

6=-4096 * -1/4^6−1 6=-4096 * -1/4^5 6=-4096 * -1/4^5 I'm stuck...

OpenStudy (anonymous):

\[\frac{ -4096 }{ 4*4*4*4*4 }\]

OpenStudy (anonymous):

\[\frac{ -4096 }{ 1024 }\] \[-4\]

OpenStudy (anonymous):

hence sixth term is -4

OpenStudy (anonymous):

Thanks, still not sure how you got it...but thanks.

OpenStudy (anonymous):

Mind helping me with a different problem?

OpenStudy (anonymous):

Given the geometric sequence where a1 = 4 and the common ratio is 3, what is the domain for n?

OpenStudy (anonymous):

i just solve the equation

OpenStudy (anonymous):

a1 = 4 and r=3 it signifies that it is an increasing gp so it can goes upto infinity

OpenStudy (kmeis002):

By definition, a sequence is defined as a collection of elements who can be ordered in accordance with the cardnality of the natural numbers. While n can be any integer, we tend to stick to the natural numbers. the domain for this sequence, for most sequences you will deal with in that class, is the natural numbers \[ n \in \mathbb{N} = \left \{1,2,3,4,5,6,7,8,9,...\right \}\]

OpenStudy (anonymous):

Would it be all integers where n ≥ 1? Or maybe all real numbers?

OpenStudy (kmeis002):

The input (domain) of a sequence cant be the real numbers, you must be careful. Saying \[n \geq 1 \] Is too vague, or just plain wrong. It must be a natural number (counting numbers)

OpenStudy (anonymous):

The answer choices are: n ≥ 1 n > 1 n ≥ 4 all real numbers That's why I thought of n ≥ 1

OpenStudy (kmeis002):

Out of those choices, the one you chose was the best, but the answers are poorly written.

OpenStudy (anonymous):

I got the previous question wrong...

OpenStudy (anonymous):

The answer was 4, not -4

OpenStudy (anonymous):

Thanks so much- like @bbbbbbbbbbbbbbb said, the correct answer was 4 :)